OpenStudy (anonymous):

What is the next prime number after 7?

OpenStudy (anonymous):

no 8 since \(8=2^3\)

OpenStudy (anonymous):

not 9 since \(9=3^2\)

OpenStudy (anonymous):

not \(10\) since \(10=2\times 5\)

OpenStudy (anonymous):

try 11

OpenStudy (anonymous):

its 11 :) i know im late XD

OpenStudy (anonymous):

i didnt i was agreeing with them..

OpenStudy (anonymous):

but ok..